OpenStudy (anonymous):

How would I solve for the real solution of 1/(x+1)^2 = 1/x+1 +2

OpenStudy (aum):

Is it \[ \Large \frac{1}{(x+1)^2} = \frac{1}{x+1} + 2 \] ? If so, multiply throughout by \(\large (x+1)^2\) and simplify and solve the quadratic equation.

OpenStudy (anonymous):

Another (but essentially identical) way to approach this is to let \(y=\dfrac{1}{x+1}\), then you have the quadratic \[\begin{align*}y^2&=y+2\\ y^2-y-2&=0\\ (y-2)(y+1)&=0\\ y&=2,~-1 \end{align*}\] Back substitute and solve for \(x\).